Please provide a detailed ‘H-NMR and C-NMR analysis for both the reactant and product. For the ‘H-NMR analysis please include:
1) on the reactant and product the types of protons and their peaks.
2) splitting pattern
3) intergration ratio.
CHO Woolt Me or Reatant Meor Product
Answer #1

Solution:- CHO Reactant Reactant Product IHNMR 13CNMR a = 55.8 b= 114.8 a = 3.83, (s, 3H) b = 7.18, (d, 2H) c=7.88, (d, 2H) d